OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –3 5 0 4 Line 2 x y 0 4 6 2 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

so did you try to solve it what you got ??

OpenStudy (anonymous):

was i suppose to subtract this divide?

Nnesha (nnesha):

okay so they gave you two table you should have two equations find slope and y -intercept of both table first :)

OpenStudy (anonymous):

was i suppose to subtract this divide?\[\frac{ y2 }{ x2}-\frac{ xy1 }{ x1 }\]

OpenStudy (anonymous):

sorry thats suppose ti be a y1

Nnesha (nnesha):

yesh first y2 -y1 and x2 -x1 and then divide

Nnesha (nnesha):

that's fine i can understand latex :P

OpenStudy (anonymous):

ya i got

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 } and \frac{ 2 }{ 6 }\]

OpenStudy (anonymous):

am i right?

Nnesha (nnesha):

first one is right

OpenStudy (anonymous):

i just don't know how to put them on the graph

Nnesha (nnesha):

second one try that again and reduce 2/6

OpenStudy (anonymous):

\[\frac{ -2 }{ 6 } sorr\]

OpenStudy (anonymous):

sorry*

Nnesha (nnesha):

yesh that's right so reduce that

OpenStudy (anonymous):

\[-\frac{ 1 }{ 3 }\]

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

now what i do

Nnesha (nnesha):

wait what was your slope for first table ??

OpenStudy (anonymous):

umm

Nnesha (nnesha):

for first one what you got -1/3 or 1/3 ??

OpenStudy (anonymous):

sorry

Nnesha (nnesha):

what is that that's a different table ??:O

OpenStudy (anonymous):

umm -1/3

OpenStudy (anonymous):

haha sorry AGAIN

OpenStudy (anonymous):

Line 1 x y -3 5 0 4 Line 2 x y 0 4 6 2

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

yesh so both table have same slope -1/3 -1/3

Nnesha (nnesha):

both have same slope mean both equation are parallel and that's mean our answer should be ??

OpenStudy (anonymous):

\[\frac{ -2 }{ }\]

Nnesha (nnesha):

what is that ??

OpenStudy (anonymous):



OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

can i just have the answer

Nnesha (nnesha):

nope sorry not at alll :(

OpenStudy (anonymous):

lol ok

Nnesha (nnesha):

you just need just one definition and then you done

OpenStudy (anonymous):

what defintion

Nnesha (nnesha):

if both lines are parallel then answer should be what ?? if 2 equations have same slope then lines are parallel

OpenStudy (anonymous):

Inconsistent

OpenStudy (anonymous):

meaning no solution

OpenStudy (anonymous):

xD

Nnesha (nnesha):

 parallel mean never touch with each ohter

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

so Inconsistent

OpenStudy (anonymous):

meaning no solution A

Nnesha (nnesha):

so yes not soulution

Nnesha (nnesha):

solution**

OpenStudy (anonymous):

thanx aare you a girl or a boy?

Nnesha (nnesha):

my pleasure :)

OpenStudy (anonymous):

I'm A girl

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

ookay.....